~~~~~~~1. The graph shows how 10,000 is invested. After 10% of the amount that is invested in bonds is reinvested in CD's , how much money is invested in CD's?
Investment Portfolio
Bonds -- 50%
Stocks -- 25%
Mutual Funds -- 15%
CD --10%
~~~~~~~9. Use the table below. Half of the students that made an A failed the last test and now have a B. What percent of the class now has a B average in math class?
Average in Math---Number of Students
A---8
B---12
C---5
D---4
E---3
My answer for the first one was 15%. For the second question, 37.5%. They were both marked wrong. Please help explain. Thank you!

Answer by Eviecker(3)

First question asked for how much money NOT what percent, so your answer should be in value/number NOT percent.
Bonds = 50% = 5000
Stocks = 25% = 2500
Mutual Funds = 15% = 1500
CD = 10% = 1000
After that 10% of the amount that is invested in BONDS is reinvested in CD's.
So take 10% out of 5000 which equals 500.
Now add 500 to 1000(the original amount invested in CD) which equals 1500.
So how much money is invested in CD's now?
Answer: 1500 is invested in CD's.
---------------------------------------------------------------------------
For the second question,
Add all numbers of students.
8 + 12 + 5 + 4 + 3 = 32
There are 32 students altogether.
Then, find the number of students that made an A but failed the last test and now have a B.
1/2 x 8 = 4
Now add 4 to 12(the original number of students who got a B) which equals 16.

Finally, use a proportion to find the percentage of the class now that has a B average in math class.
Number of students who have a B average/Total numbers of students = x/100%
16/32 = x/100%
16(100) = 32x cross multiply
1600 = 32x
1600/32 = 32x/32 divide both sides by 32
50 = x
So what percent of the class now has a B average in math class?
The answer is 50%.